# FaVoR: Features via Voxel Rendering for Camera Relocalization

### v1

@@ Abstract
Camera relocalization methods range from dense image alignment to direct camera pose regression from a query image. Among these, sparse feature matching stands out as an efficient, versatile, and generally lightweight approach with numerous applications. However, feature-based methods often struggle with significant viewpoint and appearance changes, leading to matching failures and inaccurate pose estimates. To overcome this limitation, we propose a novel approach that leverages a globally sparse yet locally dense 3D representation of 2D features. By tracking and triangulating landmarks over a sequence of frames, we construct a sparse voxel map optimized to render image patch descriptors observed during tracking. Given an initial pose estimate, we first synthesize descriptors from the voxels using volumetric rendering and then perform feature matching to estimate the camera pose. This methodology enables the generation of descriptors for unseen views, enhancing robustness to view changes. We extensively evaluate our method on the 7-Scenes and Cambridge Landmarks datasets. Our results show that our method significantly outperforms existing state-of-the-art feature representation techniques in indoor environments, achieving up to a 39% improvement in median translation error. Additionally, our approach yields comparable results to other methods for outdoor scenarios while maintaining lower memory and computational costs. Our code is open source. [^1] Code will be released upon paper acceptance.
@@ 1 Introduction
Despite extensive research, visual localization remains a significant challenge, particularly under wide viewpoint and appearance changes. Various strategies have been proposed to improve localization performance in difficult scenarios. Sequence-based place recognition, for instance, exploits consistency across frames to identify previously visited locations [27], while geometric constraints derived from point cloud alignment and 3D-3D matching can provide additional, useful information [23, 24, 47]. Ultimately, the effectiveness of many of these strategies depends upon the reliability and robustness of low-level feature matching, which is used extensively to determine the camera pose. Recent efforts to enhance visual localization have explored techniques based on neural radiance fields (NeRF) [26]. These approaches build upon dense or semi-dense simultaneous localization and mapping (SLAM) frameworks [26, 41, 33, 9, 52] by either synthesizing frames for feature point extraction or by directly performing photometric alignment. Other techniques exploit neural view synthesis to render a dense descriptor space in order to localize in a known map [28, 21], or to capture the appearance variations of a sparse set of feature points observed from a query viewpoint [14]. While these methods do improve performance, they come with some significant drawbacks. The dense descriptor representation [21, 28] often requires more training times and memory, since a dense and fully optimized radiance field must be learned prior to deployment. Furthermore, works that synthesize sparse descriptors [14] struggle to render descriptors with a large number of channels.
@@
Building on these insights, we introduce FaVoR, a new feature rendering method that exploits a pre-trained neural network to extract robust features and a sparse voxel-based formulation to encode and render feature descriptors in 3D space. Our scene representation, illustrated in Figure 1, is globally sparse but locally dense, allowing for efficient extraction of view-conditioned 3D point descriptors from any query camera pose. During training, we extract and track keypoints from a sequence of frames and triangulate 3D landmarks by leveraging known camera poses aligning our approach with existing online localization pipelines [30], which are inherently sequential. Each landmark is represented as a voxel and optimized through volumetric rendering, allowing the associated descriptor to be rendered as seen from a novel view. In turn, we are able to carry out low-level feature matching under wide viewpoint changes. Unlike prior techniques that learn dense radiance fields [28, 21], FaVoR encodes sparse landmark descriptors only, offering a favorable tradeoff between descriptor robustness and resource efficiency. FaVoR reduces the computational burden associated with more dense models, improving scalability in practical applications. Moreover, instead of using a neural network to render descriptors from different views [28, 21, 14], we employ an explicit voxel representation and trilinear interpolation for descriptor rendering, accelerating both the training and rendering processes. We extensively evaluate our approach on the 7-Scenes [40] and Cambridge Landmarks [17] datasets. Notably, FaVoR significantly outperforms existing state-of-the-art implicit feature rendering methods in the indoor 7-Scenes [40] dataset, where we demonstrate an improvement of up to 39% in median translation error.
@@
  • We demonstrate using the 7-Scenes [40] and Cambridge Landmarks [17] datasets that our system is capable of rendering descriptors faithfully from unseen views; FaVoR has improved camera pose estimation performance compared to other state-of-the-art implicit descriptors rendering techniques [14, 28, 21].
@@ 2 Related Work
We divide traditional visual localization systems into three categories according to the scene representation employed: image-based, structure-based, and hybrid. This section provides an overview of related work in these categories and then contrasts traditional methods with newer techniques (collectively called scene feature renderers) that synthesize new descriptors based on the query pose.
@@
Scene feature rendering methods (SFRMs) form a newer class of visual localization techniques that directly render 3D point descriptors from novel views in order to localize. These methods potentially overcome one of the major limitations of traditional feature extractors, that is, the lack of viewpoint invariance. Although learning-based approaches have led to improvements in feature extraction and matching, the network training process still relies on 2D images, and hence the learned descriptors are closely tied to image appearance and viewpoint. To address the problem of viewpoint invariance, FQN [14] learns to model the appearance of feature points observed from different views, but struggles to render high-dimensional descriptors (see [14, Section 6]). Moreau et al. propose CROSSFIRE in [28], a NeRF model that builds a dense scene representation in descriptor space, rather than in the RGB space, allowing for continuous descriptor rendering across different views. Most recently, Liu et al. describe NeRF-loc [21], an approach that leverages the depth information provided by RGB-D cameras to create an accurate NeRF model and then to learn 3D descriptors that can be matched with new views via a multi-headed attention mechanism. Despite the pose estimation accuracy obtained, these methods are limited by training time and memory usage, as they require a full 3D model of the environment. Consequently, they are unsuitable for large-scale and real-time localization tasks. Our approach relies on volumetric rendering but for a sparse set of voxels only, derived from features tracked over a sequence of RGB input images.
@@ 3 Methodology
We track a sparse set of landmarks through a sequence of training images. We then use volumetric rendering to optimize a local, dense voxel grid representation of each landmark, rendering the 2D descriptor patches observed during tracking, (see Figure 1). During inference we iteratively refine the camera estimate by querying the optimized voxel set to determine each landmark’s appearance (i.e., feature) based on an initial camera pose guess, as proposed by [14]. The rendered descriptors are then matched with the features in the query image, and these correspondences are used to compute the camera pose by solving the PnP problem [19] within a RANSAC scheme [12, 8]. We refer to this iterative process as Render+PnP-RANSAC. In the following subsections, we describe each step in more detail.
@@ 3.1 Landmark Tracking
Given a track, we triangulate the landmark position $\boldsymbol{\ell}_{j}$ in the world frame. An initial estimate of $\boldsymbol{\ell}_{j}$ is found using the direct linear transform algorithm [15] and refined by minimizing the reprojection error in a Levenberg–Marquardt optimization process. To account for the presence of outliers, we apply a robust cost [13] in the optimization process. To ensure the numerical stability of the optimization, we used the inverse depth parameterization of the landmark position in the camera frame. The measurement model and cost function are described in detail in our supplementary materials. Our approach is designed to integrate seamlessly into the standard front-end pipeline for robot localization, making it suitable for existing online vision-based localization systems like visual-inertial odometry (VIO) [30, 10].
@@ 3.2 Voxel Creation
We instantiate a new voxel $\mathcal{V}_{j}$, centred at $\boldsymbol{\ell}_{j}$, for each track that is longer than a certain, minimum length, as common localization system do for persistent landmarks selection [30]. The landmark associated with a track must be visible from many poses to provide useful localization information. Each voxel $\mathcal{V}_{j}$ consists of a grid, containing a number of smaller subvoxels, with a resolution of $R\times R\times R$. Each node (i.e., vertex) of the resulting grid stores a vector of size $C$, as the descriptor channels provided by $\mathcal{F}$. To determine the appropriate size of the voxel (and its subvoxels), we first compute the Euclidean distance $l_{ij}$ from $\mathbf{T}_{i}$ to the point $\boldsymbol{\ell}_{j}$, for each patch $\mathbf{P}_{ij}$. Then, given the patch size of $S\times S$ pixels, we estimate the voxel size $s_{\mathcal{V}_{j}}$ (in meters) as
@@
where $f$ is the camera focal length. We choose the minimum voxel size since this is sufficient to capture the information required to render the descriptor associated with the keypoint $\mathbf{k}_{ij}$. Additionally, each voxel $\mathcal{V}_{j}$ is associated with a density grid of the same resolution and size as the descriptors one, but with nodes of size one instead of $C$. The density grid is used in the volumetric rendering process described in Section 3.3.
@@ 3.3 Descriptor Learning and Rendering
Once the voxels are created, we can train our system to render the descriptor patches observed along the associated tracks. The training process is similar to the method described in [26], but with notable differences, such as the absence of a multi-layer perceptron (MLP) for view-dependent rendering. For all patches and poses $i\in\mathcal{S}_{j}$, we trace rays from the camera center $\mathbf{T}_{i}$ that pass through each element of the patch $\mathbf{P}_{ij}$. Each of the rays $\mathbf{r}$ intersects the voxel grid $\mathcal{V}_{j}$ and its associated density grid at two points, one closer to the camera, $\mathbf{p}_{n}$, and one farther away, $\mathbf{p}_{f}$. We take $N$ samples, $\mathbf{d}_{t}\in\mathbb{R}^{C}$ and $\hat{\sigma_{t}}\in\mathbb{R}$, with $t=1,\dots,N$, along the ray between the two intersection points, using trilinear interpolation from both $\mathcal{V}_{j}$ and the density grid, respectively. This process follows the volume rendering approach proposed in [25], but instead of rendering an RGB color, we render a feature descriptor,
@@
$\hat{\mathbf{d}}_{ij}^{uv}$ $=\sum_{t=1}^{N}T_{t}\cdot(1-\exp{(-\hat{\sigma_{t}}\delta)})\cdot\mathbf{d}_{t},$ (4)
@@
where $\hat{\mathbf{d}}_{ij}^{uv}$ is the estimated descriptor rendered for the landmark $\boldsymbol{\ell}_{j}$ seen from the pose $\mathbf{T}_{i}$ at the pixel location $(u,v)$ in patch $\mathbf{P}_{ij}$. To learn the descriptor and density grids, we want to ensure that $\hat{\mathbf{d}}_{ij}^{uv}$ in the descriptor vector space is as close as possible in norm and direction to the ground truth descriptors $\mathbf{d}_{ij}^{uv}\in\mathbf{P}_{ij}$ extracted by $\mathcal{F}$ for all the $(u,v)\in\{(0,0),(0,1),\dots,(S,S)\}$ in the patch. We use the mean squared error (MSE) and cosine similarity losses to enforce the norm and direction of the rendered descriptor, respectively, along with total variation regularization [34] to ensure a smooth representation of the patches $\mathbf{P}_{ij}$. For the density, we use the same entropy loss as described in [42]. In the supplementary material, we report on the used losses in more detail. Note that we apply the training process to each voxel independently of all the other voxels, enabling a potential parallelization of the training process.
To render the descriptors from the voxels in the scene seen from a novel view, our system requires an initial guess $\mathbf{\hat{T}}_{q}$ of the query pose $\mathbf{T}_{q}$ that we aim to estimate. We assume this estimate is provided, as is typical in robotics localization systems where a prior pose estimate is available. Given the pose $\mathbf{\hat{T}}_{q}$ and the set of voxels $\mathcal{V}=\{\mathcal{V}_{0},\mathcal{V}_{1},\dots,\mathcal{V}_{J}\}$ in the scene, we can render the descriptors for all of the landmarks that are visible from a given query pose (this rendering may include points that are occluded because of the lack of depth information). For each $\mathcal{V}_{j}$, we trace a ray from the query camera pose $\mathbf{\hat{T}}_{q}$ to the voxel-grid center $\boldsymbol{\ell}_{j}$ (i.e., the position of the landmarks). Following Equations (3), (4) and (5), we then perform volumetric rendering along the ray to obtain the expected descriptor as seen from $\mathbf{\hat{T}}_{q}$.
@@
Figure 2: Visualization of similarity response. We render a feature tracked during training using the Alike-l descriptor from an unseen view. On the left, a) displays the ground truth positions of the rendered feature points, obtained by projecting the triangulated landmarks on the camera plane, in red. At the same time, b) and c) show the similarity response between the rendered features and the target image dense descriptor map. The yellow color indicates a strong response, concentrated around the feature positions shown in a), demonstrating the effectiveness of our descriptor rendering approach. The small circle in blue is the circle center at the highest score response, the red circle is centered at the project landmark position. The top three score peaks per keypoint are 0.897, 0.868, and 0.836 for b) and 0.856, 0.786, and 0.737 for c) obtained by performing non-maxima suppression with a ray of 3 pixels, our method effectively renders descriptors that are distinctive in the image.
Once all the descriptors visible from $\mathbf{\hat{T}}_{q}$ are rendered, we can find 2D-3D correspondences with the query image $\mathbf{I}_{q}$. A feature extractor generally finds sparse 2D keypoints in the query image and associates each of them with a descriptor extracted from a dense descriptor map given by $\mathcal{F}$ [49, 11]. To match our rendered descriptors to those extracted from the query image, we find the correspondences that provide the highest similarity score above a certain threshold; i.e., we compute a similarity matrix between the two descriptor sets, and after thresholding it, we consider only the descriptor couples with the maximum similarity response. Figure 2 shows the similarity response of a rendered descriptor with the dense descriptor map extracted by the Alike-l [50] feature extractor from an unseen view.
Once all the rendered features are matched and the putative 2D-3D correspondences are available, we use the PnP [19] in a RANSAC [8] scheme to determine the pose of the camera. Our feature representation allows us to render descriptors from novel views, enabling the usage of the iterative Render+PnP-RANSAC refinement procedure. This process highlights the distinctive characteristics of the proposed 3D feature representation. As the estimated query pose $\mathbf{\hat{T}}_{q}$ converges to $\mathbf{T}_{q}$, the rendered descriptors increasingly match the query image descriptors’ appearance, resulting in a higher number of correspondences (see Section 4).

### v4

@@ Abstract
Camera relocalization methods range from dense image alignment to direct camera pose regression from a query image. Among these, sparse feature matching stands out as an efficient, versatile, and generally lightweight approach with numerous applications. However, feature-based methods often struggle with significant viewpoint and appearance changes, leading to matching failures and inaccurate pose estimates. To overcome these limitations, we propose a novel approach that leverages a globally sparse but locally dense 3D representation of 2D features. By tracking and triangulating landmarks over a sequence of frames, we construct a sparse voxel map optimized to render image patch descriptors observed during tracking. Given an initial pose estimate, we first synthesize descriptors from the voxels using volumetric rendering and then perform feature matching to estimate the camera pose. This method enables the generation of descriptors for unseen views, enhancing robustness to viewpoint changes. We evaluate our method on the 7-Scenes and Cambridge Landmarks datasets. Our results show that our approach significantly outperforms existing state-of-the-art feature representation techniques in indoor environments, achieving up to a 39% improvement in median translation error. Additionally, our approach yields comparable results to other methods for outdoor scenes but with lower computational and memory footprints.
Supplementary Material: For additional details, please visit: https://papers.starslab.ca/favor
@@ 1 Introduction
Despite extensive research, visual localization remains a significant challenge, particularly under wide viewpoint and appearance changes. Various strategies have been proposed to improve localization performance in difficult scenarios. Sequence-based place recognition, for instance, exploits consistency across frames to identify previously visited locations [27], while geometric constraints derived from point cloud alignment and 3D-3D matching can provide additional, useful information [23, 24, 47]. Ultimately, the effectiveness of many of these strategies depends upon the reliability and robustness of low-level feature matching, which is used extensively to determine the camera pose. Recent efforts to enhance visual localization performance have explored techniques based on neural radiance fields (NeRF) [26]. These approaches build upon dense or semi-dense simultaneous localization and mapping (SLAM) frameworks [26, 41, 33, 9, 52] by either synthesizing frames for feature point extraction or by directly performing photometric alignment. Other techniques exploit neural view synthesis to render a dense descriptor space in order to localize in a known map [28, 21], or to capture the appearance variations of a sparse set of feature points observed from a query viewpoint [14]. While these methods do improve performance, they come with some significant drawbacks. The dense descriptor representation [21, 28] often requires more training time and memory, since a dense and fully optimized radiance field must be learned prior to deployment. Furthermore, works that synthesize sparse descriptors [14] struggle to render descriptors with a large number of channels.
@@
Building on these insights, we introduce FaVoR, a new feature rendering method that exploits a pre-trained neural network to extract features and a sparse voxel-based formulation to encode and render feature descriptors in 3D space. Our scene representation, illustrated in Figure 1, is globally sparse but locally dense, allowing for efficient extraction of view-conditioned 3D point descriptors from any query camera pose. During training, we extract and track keypoints across a sequence of frames, then triangulate 3D landmarks using known camera poses. This approach mirrors existing online localization pipelines [30], which are inherently sequential. Each landmark is represented as a voxel and optimized through volumetric rendering, allowing the associated descriptor to be rendered from novel views. In turn, we are able to carry out low-level feature matching under wide viewpoint changes. Unlike prior techniques that learn dense radiance fields [28, 21], FaVoR encodes sparse landmark descriptors only, offering a favorable tradeoff between descriptor robustness and resource efficiency. FaVoR reduces the computational burden associated with more dense models, improving scalability in practical applications. Moreover, instead of using a neural network to render descriptors [28, 21, 14], we employ an explicit voxel representation and trilinear interpolation for rendering, accelerating both the training and rendering processes. We extensively evaluate our approach on the 7-Scenes [40] and Cambridge Landmarks [17] datasets. FaVoR significantly outperforms state-of-the-art implicit feature rendering approaches on the indoor 7-Scenes dataset [40], achieving up to a 39% reduction in median translation error, and delivers comparable performance on outdoor scenes.
@@
  • We demonstrate using the 7-Scenes [40] and Cambridge Landmarks [17] datasets that our system is capable of rendering descriptors faithfully from unseen views; FaVoR improves camera pose estimation performance compared to other state-of-the-art implicit descriptors rendering techniques [14, 28, 21].
@@ 2 Related Work
We divide traditional visual localization systems into three categories according to the scene representation employed: image-based, structure-based, and hybrid. This section provides an overview of related work in these categories and contrasts traditional methods with newer techniques (collectively called scene feature renderers) that synthesize new descriptors based on a query pose.
@@
Scene feature rendering methods (SFRMs) form a newer class of visual localization techniques that directly render 3D point descriptors from novel views in order to localize. These methods potentially overcome one of the major limitations of traditional feature extractors, that is, the lack of viewpoint invariance. Although learning-based approaches have led to improvements in feature extraction and matching, the network training process still relies on 2D images, and hence the learned descriptors are closely tied to image appearance and viewpoint. To address the problem of viewpoint invariance, FQN [14] learns to model the appearance of feature points observed from different views, but struggles to render high-dimensional descriptors (see [14, Section 6]). Moreau et al. propose CROSSFIRE in [28], a NeRF model that builds a dense scene representation in descriptor space, rather than in the RGB space, allowing for continuous descriptor rendering across different views. Most recently, Liu et al. describe NeRF-loc [21], an approach that leverages the depth information provided by an RGB-D camera to create an accurate NeRF model and then to learn 3D descriptors that can be matched with new views via a multi-headed attention mechanism. Despite the pose estimation accuracy obtained, these methods are limited by training time and memory usage, as they require a full 3D model of the environment. Consequently, they are unsuitable for large-scale and real-time localization tasks. Our approach relies on volumetric rendering but for a sparse set of voxels only, derived from features tracked over a sequence of RGB input images.
@@ 3 Methodology
We track a sparse set of landmarks through a sequence of training images. We then use volumetric rendering to optimize a local, dense voxel grid representation of each landmark, rendering the 2D descriptor patches observed during tracking (see Figure 1). At inference time, we iteratively refine the camera estimate by querying the optimized voxel set to determine each landmark’s appearance (i.e., feature) based on an initial camera pose guess, as proposed by [14]. The rendered descriptors are then matched with the features in the query image, and these correspondences are used to compute the camera pose by solving the PnP problem [19] within a RANSAC scheme [12, 8]. We refer to this iterative process as Render+PnP-RANSAC. In the following subsections, we describe each step in more detail.
@@ 3.1 Landmark Tracking
Given a track, we triangulate the landmark position $\boldsymbol{\ell}_{j}$ in the world frame. An initial estimate of $\boldsymbol{\ell}_{j}$ is found using the direct linear transform algorithm [15] and refined by minimizing the reprojection error in a Levenberg–Marquardt optimization technique. To account for the presence of outliers, we apply a robust cost in the optimization process [13]. To ensure the numerical stability of the optimization, we used the inverse depth parameterization of the landmark position in the camera frame. The measurement model and cost function are described in detail in our supplementary materials. Our approach is designed to integrate seamlessly into standard front-end pipelines for robot localization, making it suitable for existing online vision-based localization systems like visual-inertial odometry (VIO) [30, 10].
@@ 3.2 Voxel Creation
We instantiate a new voxel $\mathcal{V}_{j}$, centred at $\boldsymbol{\ell}_{j}$, for each track that is longer than a certain, minimum length, as common localization system do for persistent landmarks selection [30]. The landmark associated with a track must be visible from many poses to provide useful localization information. Each voxel $\mathcal{V}_{j}$ consists of a grid, containing a number of smaller subvoxels, with a resolution of $R\times R\times R$. Each node (i.e., vertex) of the resulting grid stores a vector of size $C$, corresponding to the descriptor channels provided by $\mathcal{F}$. To determine the appropriate size of the voxel (and its subvoxels), we first compute the Euclidean distance $l_{ij}$ from $\mathbf{T}_{i}$ to the point $\boldsymbol{\ell}_{j}$, for each patch $\mathbf{P}_{ij}$. Then, given the patch size of $S\times S$ pixels, we estimate the voxel size $s_{\mathcal{V}_{j}}$ (in metres) as
@@
where $f$ is the camera focal length. We choose the minimum voxel size since this is sufficient to capture the information required to render the descriptor associated with the keypoint $\mathbf{k}_{ij}$. Additionally, each voxel $\mathcal{V}_{j}$ is associated with a density grid of the same resolution and size as the descriptors grid, but with nodes of size one instead of $C$. The density grid is used in the volumetric rendering process described in Section 3.3.
@@ 3.3 Descriptor Learning and Rendering
Once the voxels are created, we can train our system to render the descriptor patches observed along the associated tracks. The training process is similar to the method described in [26], but with notable differences, such as the absence of a multi-layer perceptron (MLP) for view-dependent rendering. For all patches and poses $i\in\mathcal{S}_{j}$, we trace rays from the camera center $\mathbf{T}_{i}$ that pass through each element of the patch $\mathbf{P}_{ij}$. Each of the rays $\mathbf{r}$ intersects the voxel grid $\mathcal{V}_{j}$ and its associated density grid at two points, one closer to the camera, $\mathbf{p}_{n}$, and one farther away, $\mathbf{p}_{f}$. We take $N$ samples, $\mathbf{d}_{t}\in\mathbb{R}^{C}$ and $\hat{\sigma_{t}}\in\mathbb{R}$, with $t=1,\dots,N$, along the ray between the two intersection points, using trilinear interpolation from both $\mathcal{V}_{j}$ and the density grid, respectively (see Figure 1).This process follows the volume rendering approach proposed in [25], but instead of rendering an RGB color, we render a feature descriptor,
@@
$\hat{\mathbf{d}}_{ij}^{uv}$ $=\sum_{t=1}^{N}T_{t}\left(1-\exp{(-\hat{\sigma_{t}}\delta)}\right)\,\mathbf{d}_{t},$ (4)
@@
where $\hat{\mathbf{d}}_{ij}^{uv}$ is the estimated descriptor for landmark $\boldsymbol{\ell}_{j}$ seen from the pose $\mathbf{T}_{i}$ at the pixel location $(u,v)$ in patch $\mathbf{P}_{ij}$. To learn the descriptor and density grids, we want to ensure that $\hat{\mathbf{d}}_{ij}^{uv}$, in the descriptor vector space, is as close as possible in norm and direction to the ground truth descriptors $\mathbf{d}_{ij}^{uv}\in\mathbf{P}_{ij}$ extracted by $\mathcal{F}$ for all the $(u,v)\in\{(0,0),(0,1),\dots,(S,S)\}$ in the patch. We use the mean squared error (MSE) and cosine similarity losses to drive the rendered descriptor’s norm and direction as close as possible to the target feature, along with total variation regularization [34] to ensure a smooth representation of the patches $\mathbf{P}_{ij}$. For the density, we use the same entropy loss as described in [42]. In our supplementary material, we provide more details on the losses used. Note that we apply the training process to each voxel independently of all the other voxels, enabling parallelization of the training process.
To render the descriptors from the voxels in the scene observed from a novel view, our system requires an initial guess $\mathbf{\hat{T}}_{q}$ for the query pose $\mathbf{T}_{q}$ that we aim to estimate. We assume that this estimate is provided, as is typical in robotics localization systems where a prior pose estimate is available. Given the pose $\mathbf{\hat{T}}_{q}$ and the set of voxels $\mathcal{V}=\{\mathcal{V}_{0},\mathcal{V}_{1},\dots,\mathcal{V}_{J}\}$ in the scene, we can render the descriptors for all of the landmarks that are visible from a given query pose (this rendering may include points that are occluded because of the lack of depth information). For each $\mathcal{V}_{j}$, we trace a ray from the query camera pose $\mathbf{\hat{T}}_{q}$ to the voxel-grid center $\boldsymbol{\ell}_{j}$ (i.e., the position of the landmarks). Following Equations (3), (4) and (5), we then perform volumetric rendering along the ray to obtain the expected descriptor as seen from $\mathbf{\hat{T}}_{q}$.
@@
Figure 2: Visualization of similarity response. We render a feature tracked during training using the Alike-l descriptor from an unseen view. On the left, a) displays the ground truth positions of the rendered feature points, obtained by projecting the triangulated landmarks on the camera plane, in red. At the same time, b) and c) show the similarity response between the rendered features and the target image dense descriptor map. The yellow colour indicates a strong response, concentrated around the feature positions shown in a), demonstrating the effectiveness of our descriptor rendering approach. The small circle in blue is the circle centre at the strongest response, the red circle is centred at the project landmark position. The top three response peaks per keypoint are 0.897, 0.868, and 0.836 for b) and 0.856, 0.786, and 0.737 for c) obtained by performing non-maxima suppression with a ray of three pixels, our method effectively renders descriptors that are distinctive in the image.
Once all the descriptors visible from $\mathbf{\hat{T}}_{q}$ are rendered, we can find 2D-3D correspondences with the query image $\mathbf{I}_{q}$. A feature extractor generally finds sparse 2D keypoints in the query image and associates each of them with a descriptor extracted from a dense descriptor map given by $\mathcal{F}$ [49, 11]. To match our rendered descriptors to those extracted from the query image, we find the correspondences that provide the highest similarity score, above a certain threshold; that is, we compute a similarity matrix between the two descriptor sets, and after thresholding, we consider only the descriptor pairs with the maximum similarity response. Figure 2 shows the similarity response of a rendered descriptor with the dense descriptor map extracted by the Alike-l [50] feature extractor from an unseen view.
Once all the rendered features are matched and the putative 2D-3D correspondences are available, we use PnP [19] in a RANSAC [8] scheme to determine the camera pose. Our feature representation allows us to render descriptors from novel views, enabling the usage of the iterative Render+PnP-RANSAC refinement procedure. This process highlights the distinctive characteristics of the proposed 3D feature representation. As the estimated query pose $\mathbf{\hat{T}}_{q}$ converges to $\mathbf{T}_{q}$, the rendered descriptors increasingly match the query image descriptors’ appearance, resulting in a higher number of correspondences (see Section 4).